OpenStudy (anonymous):

I need a factoring reminder; factor. x^2+9x+20

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

what grade you in?

OpenStudy (anonymous):

What times what = x^2?

OpenStudy (anonymous):

x, right? So we have: (x + ?)(x + ?) Now we need two numbers that multiply to 20 and add to 9.

OpenStudy (anonymous):

What two numbers are those? @Cassandra_Lea_96

OpenStudy (luigi0210):

Do you mean factoring remainder as in The Factor Theorem?

OpenStudy (crissy15):

5 and 4 right @iGreen.

OpenStudy (anonymous):

sorry my computer wasn't loading but yes x*x = x^2 and 5*4 =20

OpenStudy (anonymous):

@Luigi0210 I believe so it just gave me the equation and said factor and I forgot the steps to work it out =o

OpenStudy (anonymous):

@CallMeEmmie senior

OpenStudy (crissy15):

I'm a freshman though I'm suppose to be a sophomore and if skip a grade junior if I was in public school

OpenStudy (luigi0210):

Whoops, I read the question wrong.. but yea you got it right :P

OpenStudy (anonymous):

lol it's ok =) thank you =)

OpenStudy (anonymous):

Yes, so we have: (x + 5)(x + 4)